OpenStudy (anonymous):

Central Clinic recently purchased a new computer system for a total cost of $64,735. The estimated life of the system is 6 years. The trade-in value of the system after 6 years is estimated to be $900. Using the straight-line method, what is the annual depreciation to the nearest cent?

OpenStudy (chosenmatt):

welocme to openstudy

OpenStudy (danjs):

the slope, you are given 2 points

OpenStudy (danjs):

(time, value) (0 , 64735) (6 , 900)

OpenStudy (anonymous):

Annual Depreciation =Original Cost– Salvage Value Estimated Life is the formula it gives...

OpenStudy (anonymous):

I already tried but still got it wrong

OpenStudy (danjs):

\[\frac{ (900 - 64735) }{ 6 - 0 } = -10639.17 \frac{ dollars }{ year }\]

OpenStudy (anonymous):

Thank you!

OpenStudy (danjs):

u understand?

OpenStudy (anonymous):

Well the first answer I got was 10639.1666... So i rounded the .16 to 10639.27